# The Tidal Torque Theory Revisited. I. Protohalo Angular Momentum

### v2

@@ 4 Protohalo angular momentum
Once we know the Lagrangian inertia tensor I of the protohalo (Sec. 2) and the Lagrangian tidal tensor ${\bf T_{\rm d}}$ caused by the main source of the tidal torque (Sec. 3), we can proceed to calculate the Lagrangian AM, $J^{\rm L}$, of the protohalo. That is simply the tensor product $\epsilon_{ijk}{\bf T}_{{\rm n}jl}{\bf I}_{lk}$ in any given Cartesian coordinate system. Of course, the principal semi-axes of those tensors are not aligned in general with each other (in fact, that particular configuration would lead to a null tidal torque), but they are rotated with respect to that coordinate system according to the suited rotation (or direction cosine) matrices R and ${\bf R}_{\rm d}$, functions of the Euler angles $\{\alpha,\beta,\kappa\}$ and $\{\alpha_{\rm d},\beta_{\rm d},\kappa_{\rm d}\}$, respectively. Thus, we have in general
$J_{{\rm n}i}^{\rm L}=-\epsilon_{ijk}[{\bf R_{\rm d}T}_{\rm d}{\bf R_{\rm d}^{\!T}}]_{jl}[{\bf RIR^{\!T}}]_{lk}\,,$ (40)
where I and ${\bf T_{\rm d}}$ are the tensors in the form (20) and (39), and index T denotes transpose.
Equation (40) presumes one particular configuration of the system. To find the typical Lagrangian AM of the seeds of haloes with $M$ at $t$ we must average the absolute magnitude of the components $J_{{\rm n}i}^{\rm L}$ over all possible configurations of the ‘protohalo-main source of tidal torque’ system for the joint PDF of all its properties,
@@
with $C_{\rm d}=\{e_{\rm d},p_{\rm d},\alpha_{\rm d},\beta_{\rm d},\kappa_{\rm d}\}$ and $C=\{e,p,\alpha,\beta,\kappa\}$, and the PDFs $P_{\rm pk}$ and $P^{\rm st}$ given by equations (16) and (35), respectively.
Actually, we are interested in the average of the $J_{\rm d}$ modulus rather than the average of the absolute magnitude of the components $J_{{\rm n}i}$. However, by isotropy of the universe, the averaged absolute magnitude of the three components, $|J_{{\rm n}k}|$, must be the same, implying $\langle J_{\rm d}\rangle^{2}=3\langle|J_{{\rm n}i}|\rangle^{2}$. Consequently, to calculate the average modulus we may simply calculate the average absolute magnitude of any of its components. Moreover, since for any configuration of the two objects there is another one yielding the same AM but with opposite sign, the average of $|J_{{\rm n}i}|$ over the whole composite angular surface, $(8\pi^{2})^{2}$ strad2, is the same as the average of $J_{{\rm n}i}$ over half that angular surface. We can thus average expression (40), say, over $\alpha$, $\beta$ and $\kappa$ in $8\pi^{2}$ strad2 and over $\alpha_{\rm d}$, $\beta_{\rm d}$ and $\gamma_{\rm d}$ in $28\pi^{2}/$ 2 strad2.
By doing this, we arrive at the following approximate analytic expression for the averaged Lagrangian AM (see App. B)
$J_{\rm d}^{\rm L}\!\approx\!\frac{2G\bar{\rho}^{1/3}_{0}\delta M^{5/3}}{5\times 3^{7/6}D(t_{\rm i})}\!\left(\frac{r}{R^{\rm th}}\right)^{\!\!m}\!\!H\!\left(e_{\rm d},p_{\rm d}\right)\!(1\!-\!2e)^{2/3}p^{1/3},$ (42)
where $\bar{\rho}_{0}$ is the current cosmic mean density and $H(e_{\rm d},p_{\rm d})$ is defined in equation (64).
The next step is to average over the ellipticity and prolateness of the two peaks for the corresponding probability functions $P_{\rm ep}$ (eq. [12]). The result to leading order for $M_{\rm vir}$ masses is (see App. C)
$J^{\rm L}_{\rm d}\approx 3.3\times 10^{-3}\,\frac{G\bar{\rho}^{1/3}_{0}\delta\,[g(\gamma,\gamma\nu)M]^{5/3}}{D(t_{\rm i})}\left(\frac{r}{R^{\rm th}}\right)^{\widetilde{m}},$ (43)
where $\widetilde{m}=13m/3$ and
$g(\gamma,\gamma\nu)\equiv\frac{(\gamma\nu)^{2}}{\langle x\rangle^{2}+6/5}.$ (44)
Lastly, to obtain the mean Lagrangian AM, $J^{\rm L}$, of protohaloes collapsing into haloes with $M$ at $t$ we must average $J_{\rm d}^{\rm L}$ given by equation (43) over $r$,
@@
where $P(\delta,R,r)$ is the integral over $C_{\rm d}$ and $C$ of $P(C_{\rm d},C|\delta,R,r)$ (eq. [41]), with $P_{\rm pk}=1$ and $P^{st}$ with no arguments $C_{\rm d}$ and $C$, and $r_{\rm st}(\nu,R)$ ($>R^{\rm th}$) is the radius of the sphere centred at the c.o.m. of the protohalo encompassing the nearest dominant tidal torque source (peak or hole), solution of the implicit equation
@@
being ${\cal N}_{\rm pk}(\nu,R)$ the integral of ${\cal N}_{\rm pk}(\nu,C,R)$ (eq. 15) over $C$, which leads to
@@
The typical $r_{\rm st}/R^{\rm th}$ value ($\approx 2r_{\rm st}/R$ for virial masses) as a function of $M$ is shown in Figure 1. Note that $r_{\rm st}/R^{\rm th}$ is not only constant in linear regime where $r_{\rm st}$ and $R^{\rm th}$ are proportional to $a(t)$, but it is also independent of the halo collapse time because $\nu$ (with $R^{\rm th}$ equal to the protohalo radius, different from the radius of the final virialized object) is time-invariant.
We remark that, in the case of peaks on a background $\delta_{\rm m}$ at a scale $R_{\rm m}>3R$, the same procedure followed in Salvador-Solé & Manrique (2024) would lead to a conditional number density ${\cal N}_{\rm pk}$ of such peaks that differs from that appearing in equation (46) by an additional term (to first order in $\delta_{\rm m}$) proportional to $\nu\delta_{\rm m}/\sigma_{0}$, i.e. of the same amount but opposite sign (through $\nu$) for peaks and holes. Consequently, the change in the number density of both kinds of objects would balance and $r_{\rm st}/R^{\rm th}$ would be the same. This collateral result is used in Salvador-Solé, Manrique & Agulló (2024).
Figure 1: Typical separation between the protohalo and main mass fluctuation responsible for the tidal torque, scaled to the protohalo radius, as a function of the virial mass of the final halo (independent of the redshift, for virial masses) in the Planck14 cosmology.
After performing the integral (45) with such a value of $r_{\rm st}$, we arrive at the following mean Lagrangian protohalo AM
$J^{\rm L}\approx\frac{4.95\times 10^{-3}}{\widetilde{m}+3}G\bar{\rho}^{1/3}_{0}\!\left(\frac{r_{\rm st}}{R^{\rm th}}\right)^{\widetilde{m}}[g(\gamma,\gamma\nu)M]^{5/3}\frac{\delta}{D(t_{\rm i})}.$ (48)
In Figure 2 we plot $J^{\rm L}$ as a function of the virial mass $M$ of halos (regardless of $t_{0}$).
Figure 2: Mean Lagrangian protohalo AM as a function of the virial mass of the final halo in the Planck14 cosmology.
Similarly, by integrating $J^{\rm L}$ (eq. [45]) from $R^{\rm th}$ to the median separation $r_{\rm med}$ between the protohalo and the nearest main tidal torque source, solution of the equation
$0.5=\frac{\int_{0}^{r_{\rm med}}{\rm d}r\,r^{\widetilde{m}+2}}{\int_{0}^{r_{\rm st}}{\rm d}r\,r^{\widetilde{m}+2}}=\left(\frac{r_{\rm med}}{r_{\rm st}}\right)^{\widetilde{m}+3},$ (49)
implying $r_{\rm med}=r_{\rm st}/2^{1/(\widetilde{m}+3)}$, we are led to the median Lagrangian protohalo AM, $J^{\rm L}_{\rm med}=J^{\rm L}/2$.
@@ 5 Summary and Discussion
We have applied the TTT to derive the Lagrangian AM of protohaloes centred in Gaussian triaxial peaks in CDM cosmologies, accounting for their accurate extension (mass) and ellipsoidal collapse time. These improvements have been possible thanks to the use of the CUSP formalism, making the link between haloes with $M$ at $t$ and Gaussian peaks with density contrast $\delta$ on scale $R$ in the smoothed linear Gaussian random density field at an initial time $t_{\rm i}$.
@@
The approach we have followed is innovative in the sense that not only is the protohalo a peak, but also the mass excess or defect responsible for the main tidal torque is a peak or a hole. Specifically, this mass fluctuation is found to be the largest peak or hole fitting in the neighbourhood of the protohalo (i.e. with top-hat scale $R^{\rm th}_{\rm d}$ equal to the separation $r_{\rm st}$) over the continuous $\delta(R)$ peak (or hole) trajectory starting as a twin (positive or negative) peak to that associated with the protohalo. Of course, there are many more positive and negative density fluctuations outside that trajectory in the neighbourhood of the protohalo that cause small additional tidal torques. However, these perturbations are very numerous and typically half of them slightly boosts the tidal torque caused by the main source and the other half mitigates it as they point to the opposite direction, so their global effect basically cancels.
@@
The predicted mean Lagrangian AM of protohaloes with virial mass $M$ collapsing at $t_{\rm c}$, $J^{\rm L}(M,t_{\rm c})$ is analytic, so it can be readily checked against simulations and implemented in analytic models of structure formation. Its expression is
$J^{\rm L}(M,t_{\rm c})\approx\frac{5.0\times 10^{-3}}{\widetilde{m}+3}G\bar{\rho}^{1/3}_{0}s^{\widetilde{m}}(M)\widetilde{M}^{5/3}\frac{\delta(t_{\rm c},t_{\rm i})}{D(t_{\rm i})},$ (50)
@@
where $\widetilde{m}=-13[(n+3)/2]/3\approx-2.71$ (for the CDM spectral index $n\approx-1.75$ corresponding to galactic haloes), $\bar{\rho}_{0}$ is the present mean cosmic density, $s=r_{\rm st}/R^{\rm th}$ is the time-invariant typical separation $r_{\rm s}$ between the c.o.m. of the protohalo and the main tidal torque source scaled to the protohalo size, (eq. [47]), $\widetilde{M}\equiv g(\gamma,\gamma\nu)M$, where $g$ is given by equation (44), and $\delta(t_{\rm c},t_{\rm i})$, proportional to the cosmic growth factor at $D(t_{\rm i})$ (eq. [2]), is the density contrast of Gaussian peaks at $t_{\rm i}$ that collapse ellipsoidally at $t_{\rm c}$. On the other hand, the median Lagrangian protohalo AM is $J^{\rm L}(M,t_{\rm c})/2$. Both Lagrangian AMs are, of course, independent of the (arbitrary) initial time $t_{\rm i}$ and of the cosmic time $t$ as well, their Eulerian counterparts being multiplied by $a^{2}(t)\dot{D}(t)$.
On the contrary, $J^{\rm L}$ depends on $t_{\rm c}$ through $\delta(t_{\rm c},t_{\rm i})/D(t_{\rm i})$, and on the $M$ through the factor $M^{5/3}$, as found in simulations (e.g. Sugerman, Summers & Kamionkowski 2000) apart from the slight dependence arising from typical separation between the protohalo and the main source of the tidal source, $s(M)$, which stays around $4.5$ except at the high-mass end where the halo number density rapidly falls off and $s(M)$ goes to infinity, and from the very smooth factor $g(\gamma,\gamma\nu)$ arising from the triaxial shape and relative orientation of both objects.
No need to say that the Lagrangian protohalo AM derived here only holds in linear (and moderately non-linear) regime. However, in Paper II we use this result to infer the typical AM and spin of virialised haloes as well as their specific AM profile with unprecedented accuracy.
@@ Appendix A Tidal Tensor
${\bf T}_{{\rm n}ij}\!\equiv\!\frac{a^{3}(t_{\rm i})}{D(t_{\rm i})}\frac{\partial^{2}\Phi_{\rm d}}{\partial x_{j}\partial x_{i}}\!=\!\frac{a^{3}(t_{\rm i})}{D(t_{\rm i})}\frac{4\pi\Delta[S({\bf 0})]}{\Delta[S({\bf x})]}G\delta_{\rm d}\bar{\rho}(t_{\rm i})\!\!\left(\!\frac{x_{i}x_{j}}{[a_{{\rm n}i}^{2}\!+\!S({\bf x})][a_{{\rm n}j}^{2}\!+\!S({\bf x})]}\!\left\{\sum_{k=1}^{3}\!\frac{x_{k}^{2}}{[a_{{\rm n}k}^{2}+S(\bf x)]^{2}}\!\right\}^{\!\!-1}\!\!\!\!-\delta_{ij}\!\int_{S({\bf x})}^{\infty}\!\!\frac{\Delta[S({\bf x})]\,{\rm d}s}{\Delta(s)(a^{2}_{{\rm n}i}\!+\!s)}\!\right)\!,$ (51)
@@
Since $R_{\rm d}$ is limited by the condition $R_{\rm d}^{\rm th}=r$, the c.o.m. of the protohalo is at the edge of the ellipsoidal mass excess, i.e. $\sum_{k=1}^{3}x_{k}^{2}/a^{2}_{{\rm n}k}=1$ and $S({\bf x})=0$. Consequently, the tidal tensor is simply
${\bf T}_{{\rm n}ij}=\frac{a^{3}(t_{\rm i})}{D(t_{\rm i})}4\pi G\delta_{\rm d}\bar{\rho}(t_{\rm i})\left[\frac{x_{i}x_{j}}{a^{2}_{{\rm n}i}a^{2}_{{\rm n}j}}\left(\sum_{k=1}^{3}\frac{x_{k}^{2}}{a_{{\rm n}k}^{4}}\right)^{\!-1}-\delta_{ij}U(a^{2}_{{\rm n}k})\right]\!,$ (52)
@@
$U(a^{2}_{{\rm n}k})\equiv\frac{1}{2}\int_{0}^{\infty}\frac{{\rm d}s/a^{2}_{{\rm n}j}}{1+s/a_{{\rm n}j}^{2}}\left[\prod_{k=1}^{3}\left(1+\frac{s}{a^{2}_{{\rm n}k}}\right)\right]^{-1/2}=\frac{1}{2}\int_{1}^{\infty}\,\frac{{\rm d}y}{y^{3/2}}\left[\sum_{k=1}^{3}\frac{a_{{\rm n}j}+s(y)^{2}}{a_{{\rm n}k}^{2}+s(y)}\right]^{-1}.$ (53)
Taking into account that the ellipticity and prolateness of peaks is moderate (BBKS), we have $U(a^{2}_{{\rm n}k})\approx\frac{1}{6}\int_{1}^{\infty}\,\frac{{\rm d}y}{y^{3/2}}=\frac{1}{3}$. In addition, we can neglect the terms in $1-a^{2}_{{\rm n}k}/a^{2}_{\rm d}$, where $a_{\rm d}^{2}\equiv 1/3\sum_{k=1}^{3}a^{2}_{{\rm n}k}$, in the Taylor expansion of $\sum_{k=1}^{3}x_{k}^{2}/a^{4}_{{\rm n}k}$ at $\sum_{k=1}^{3}x_{k}^{2}/(a^{2}_{{\rm n}k}a^{2}_{\rm d})$,
$\sum_{k=1}^{3}\frac{x_{k}^{2}}{a_{{\rm n}k}^{4}}=\sum_{k=1}^{3}\frac{x_{k}^{2}}{a_{{\rm n}k}^{2}a_{\rm d}^{2}}\left[1+\left(1-\frac{a_{{\rm n}k}^{2}}{a^{2}_{\rm d}}\right)+...\right]\approx\frac{1}{a_{\rm d}^{2}}\sum_{k=1}^{3}\frac{x_{k}^{2}}{a_{{\rm n}k}^{2}}=\frac{1}{a_{\rm d}^{2}}.$ (54)
Thus we arrive at the almost fully accurate expression
${\bf T}_{{\rm n}ij}=\frac{a^{3}(t_{\rm i})}{D(t_{\rm i})}\frac{4\pi}{3}G\delta_{\rm d}\bar{\rho}(t_{\rm i})\left[3\frac{x_{i}x_{j}}{a^{2}_{{\rm n}i}a^{2}_{{\rm n}j}}a_{\rm d}^{2}-\delta_{ij}\right].$ (55)
Note that in the limit $a_{{\rm n}k}=R^{\rm th}_{\rm d}$, where the two approximations above hold exactly, expression (55) recovers the Hessian of the gravitational potential at the edge ($r=R_{\rm d}^{\rm th}$) of a spherically symmetric system, as expected.

### v5

@@ 4 Protohalo angular momentum
Once we have determined the Lagrangian protohalo inertia tensor I (Sec. 2) and tidal tensor ${\bf T_{\rm d}}$ due to a dominant peak or hole (Sec. 3), we can proceed to calculate the Lagrangian AM, $J^{\rm L}$. That is simply the tensor product of those two tensors, equal to $\epsilon_{ijk}{\bf T}_{{\rm d}jl}{\bf I}_{lk}$ in a given Cartesian coordinate system. Of course, the principal axes of those tensors are not necessarily aligned with each other (if they are the tidal torque is null), but they show different spatial rotations with respect to the coordinate system of reference. Thus, we have
$J_{{\rm d}i}^{\rm L}=-\epsilon_{ijk}[{\bf R_{\rm d}T}_{\rm d}{\bf R_{\rm d}^{\!T}}]_{jl}[{\bf RIR^{\!T}}]_{lk}\,,$ (39)
where I and ${\bf T_{\rm d}}$ are the tensors oriented along their own principal axes, in the form (20) and (38), respectively, and ${\bf R}$ and ${\bf R}_{\rm d}$ are the respective rotation (or direction cosine) matrices, functions of the Euler angles $\{\alpha,\beta,\kappa\}$ and $\{\alpha_{\rm d},\beta_{\rm d},\kappa_{\rm d}\}$, index T denoting transpose.
Equation (39) is for one particular spatial configuration of the system. To find the typical Lagrangian protohalo AM we must average the absolute magnitude of components $J_{{\rm d}i}^{\rm L}$ over all possible configurations of the ‘protohalo-main tidal torque source’ system for the joint PDF of all its properties,
@@
with $C_{\rm d}=\{e_{\rm d},p_{\rm d},\alpha_{\rm d},\beta_{\rm d},\kappa_{\rm d}\}$ and $C=\{e,p,\alpha,\beta,\kappa\}$, being the PDFs $P_{\rm pk}$ and $P^{\rm st}$ given by equations (16) and (34), respectively.
Actually, what we are interested in is the average of the $J_{\rm d}$ modulus rather than the average of the absolute magnitude of the components $J_{{\rm d}i}$. However, by isotropy of the universe, the averaged absolute magnitude of any component $|J_{{\rm d}i}|$ must be the same, implying $\langle J_{\rm d}\rangle^{2}=3\langle|J_{{\rm d}i}|\rangle^{2}$. Consequently, the average modulus can be calculated from the average absolute magnitude of any given component. Moreover, since for any spatial configuration of the two objects, there is another one yielding an AM with the same modulus and opposite sign, the average of $|J_{{\rm d}i}|$ over the whole composite angular surface, $(8\pi^{2})^{2}$ strad2, is the same as the average of $J_{{\rm d}i}$ over half that angular surface. We can thus average expression (39), say, over $\alpha$, $\beta$ and $\kappa$ in $8\pi^{2}$ strad2 and over $\alpha_{\rm d}$, $\beta_{\rm d}$ and $\gamma_{\rm d}$ in $28\pi^{2}/$ 2 strad2.
By doing this (see App. B), we arrive at the following approximate analytic expression for the average Lagrangian protohalo AM
$J_{\rm d}^{\rm L}\!\approx\!\frac{2G\bar{\rho}^{1/3}_{0}\delta M^{5/3}}{5\times 3^{1/2}D(t_{\rm i})}\!\left(\frac{r}{R^{\rm th}}\right)^{\!\!m}\!\!H\!\left(e_{\rm d},p_{\rm d}\right)\!\left(\frac{1\!-\!2e}{1\!+\!e}\right)^{2/3}p,$ (41)
where $\bar{\rho}_{0}$ is the current cosmic mean density and $H(e_{\rm d},p_{\rm d})$ is a function of the ellipticity and prolateness of the peak or hole marking the tidal torque source, defined in equation (64).
Next step is to average over the ellipticity and prolateness of the two objects for the corresponding PDFs $P_{\rm ep}$ (eq. [12]). The result for virial masses to leading order is (see App. C)
$J^{\rm L}_{\rm d}\approx 0.347\,\frac{G\bar{\rho}^{1/3}_{0}\delta\,g(\gamma,\gamma\nu)M^{5/3}}{D(t_{\rm i})}\left(\frac{r}{R^{\rm th}}\right)^{5m/3},$ (42)
where
$g(\gamma,\gamma\nu)\equiv\left[\frac{(\gamma\nu)^{1/3}}{\langle x\rangle^{2}+6/5}\right]^{2}.$ (43)
Since $m=-(n+3)/2$ is negative, the peak or hole marking the mass excess or default causing the strongest tidal torque on the protohalo is that lying the closest to the protohalo c.o.m. for which $r$ is the smallest. Therefore, the mean Lagrangian AM, $J^{\rm L}$, of protohaloes collapsing at $t_{\rm c}$ in haloes with $M$ is the average of $J_{\rm d}^{\rm L}$ given by equation (42) over the separation $r$ inside the sphere of radius $r_{\rm st}$ centred at the c.o.m. of the protohalo and statistically harbouring only one such dominant peaks or holes,
@@
where, after the previous averages, $P(\delta,R,r)$ is given by equation (40), with $P_{\rm pk}=1$ and $P^{st}$ with no arguments $C_{\rm d}$ and $C$, and, as mentioned, $r_{\rm st}(\delta,R)$ ($>R^{\rm th}$) is the solution of the implicit equation
@@
where ${\cal N}_{\rm pk}(\nu,R)$ results from marginalising ${\cal N}_{\rm pk}(\nu,C,R)$ (eq. 15) with respect to $C$, which leads to
@@
The typical $r_{\rm st}/R^{\rm th}$ value ($\approx 2r_{\rm st}/R$ for virial masses) as a function of $M$ is shown in Figure 1. Note that $r_{\rm st}/R^{\rm th}$ is not only constant in linear regime where $r_{\rm st}$ and $R^{\rm th}$ are proportional to $a(t)$, but it is also independent of the halo collapse time because $\nu$ is time-invariant and weakly dependent on $M$, being $r_{\rm st}/R^{\rm th}\sim 4.5$ for all halo masses except at the very high-mass end where the halo number density rapidly falls off and $s$ goes to infinity ($s^{5m/3}$ vanishes).
Figure 1: Typical separation (independent of the redshift) between the protohalo and main mass fluctuation responsible for the tidal torque, scaled to the protohalo radius, as a function of the virial mass of current haloes in the Planck14 cosmology.
Figure 2: Mean Lagrangian protohalo AM as a function of the virial mass of present haloes in the Planck14 cosmology (solid red line). The straight dashed blue line shows the best fit to a simple $M^{5/3}$ law.
It is worth noting that, in the case of peaks conditioned to lie in a background $\delta_{\rm m}$ at a scale $R_{\rm m}>3R$, the same procedure followed in Salvador-Solé & Manrique (2024) would lead to a conditional peak number density ${\cal N}_{\rm pk}$ that differs from that appearing in equation (45) by an additional term (to first order in $\delta_{\rm m}$) proportional to $\nu\delta_{\rm m}/\sigma_{0}$, i.e. of the same amount but opposite sign (through $\nu$) for peaks and holes. Consequently, the change in the number density of both kinds of objects would balance and $r_{\rm st}/R^{\rm th}$ would remain the same. This collateral result is used in Salvador-Solé, Manrique & Agulló (2024).
After performing the integral (44), we arrive at the following mean Lagrangian protohalo AM
$J^{\rm L}\approx\frac{0.521}{5m/3+3}G\bar{\rho}^{1/3}_{0}\!\left(\frac{r_{\rm st}}{R^{\rm th}}\right)^{\frac{5m}{3}}\!g(\gamma,\gamma\nu)M^{5/3}\frac{\delta}{D(t_{\rm i})},$ (47)
plotted in Figure 2 as a function of the virial mass $M$ of haloes at $t_{\rm c}$ equal to the present time. We see that the dependence on $M$ of factors $(r_{\rm st}/R^{\rm th})^{5m/3}$ and $g(\gamma,\gamma\nu)$ is very week so that $J^{\rm L}$ is essentially proportional to $M^{5/3}$.
Similarly, by integrating $J^{\rm L}$ (eq. [42]) from $R^{\rm th}$ to the median separation $r_{\rm med}$ between the protohalo and the nearest dominant tidal torque source, solution of the equation
$0.5=\frac{\int_{0}^{r_{\rm med}}{\rm d}r\,r^{\frac{5m}{3}+2}}{\int_{0}^{r_{\rm st}}{\rm d}r\,r^{\frac{5m}{3}+2}}=\left(\frac{r_{\rm med}}{r_{\rm st}}\right)^{\frac{5m}{3}+3}$ (48)
implying $r_{\rm med}=r_{\rm st}/2^{1/(5m/3+3)}$, we are led to the median Lagrangian protohalo AM, $J^{\rm L}_{\rm med}=J^{\rm L}/2$.
@@ 5 Summary and Discussion
Using the TTT, we have derived the Lagrangian protohalo AM in CDM cosmologies within the framework of the peak model of structure formation where haloes are believed to form from the collapse of patches around triaxial peaks in the Gaussian-smoothed Gaussian random density field at an (arbitrary) initial time $t_{\rm i}$. This has been done by accounting for the first time for the accurate extension (mass) and ellipsoidal collapse time of such patches. These improvements have been possible thanks to the use of the CUSP formalism making the link between haloes with $M$ at the cosmic time $t_{\rm c}$ and Gaussian peaks with density contrast $\delta$ on scale $R$.
@@
The approach we have followed is innovative in the sense that not only are protohaloes identified to peaks, but the mass excesses or defaults responsible for the torques are also identified to peaks or holes. Specifically, we have found that, among all it nested (positive or negative) density fluctuations of different scales around a protohalo, traced by continuous $\delta(R)$ peak or hole trajectories, the dominant ones are the largest ones fitting in the neighbourhood of the protohalo (i.e. with top-hat scale $R^{\rm th}_{\rm d}$ equal to the separation $r$) whose peak at scale $R$ is twin (possibly except for the sign) to the peak marking the protohalo itself. Those ‘dominant peaks’ are indeed the most probable ones due to the correlation between peaks (or between peaks and holes) of the same scale (e.g. Salvador-Solé & Manrique 2024). And, among all those dominant tidal torque sources, the strongest tidal torque is due to that lying the closest to the protohalo. Of course, there are many more sets of nested (positive or negative) density fluctuations in the neighbourhood of the protohalo that cause small additional tidal torques. But these additional density fluctuations are very numerous and typically half of them slightly boost the tidal torque of the main source and the other half mitigates it (their tidal torque point to the opposite direction), so their global effect basically cancels.
@@
Another important accomplishment has been that the predicted mean Lagrangian AM $J^{\rm L}(M,t_{\rm c})$ of protohaloes with virial mass $M$ collapsing at $t_{\rm c}$ has the simple analytic expression
$J^{\rm L}\approx\frac{0.521}{5m/3+3}G\bar{\rho}^{1/3}_{0}s^{\frac{5m}{3}}g(\gamma,\gamma\nu)\,M^{5/3}\frac{\delta}{D(t_{\rm i})},$ (49)
@@
so it can be readily checked against simulations and implemented in analytic models of structure formation. In equation (47) $m$ is $-(n+3)/2$ in terms of the effective power index $n\approx-1.75$ of the CDM spectrum at galactic halo masses, $\bar{\rho}_{0}$ is the present mean cosmic density, $s\equiv r_{\rm st}/R^{\rm th}$ is the time-invariant and approximately mass-independent ($r_{\rm st}/R^{\rm th}\approx 4.5$) typical separation between the c.o.m. of the protohalo and the main tidal torque source scaled to the protohalo size, factor $g$ is also a very weak function of $M$ given by equation (43) and $\delta(t_{\rm c},t_{\rm i})\propto D(t_{\rm i})$ (eq. [2]) is the density contrast of Gaussian peaks at $t_{\rm i}$ that undergo ellipsoidal collapse at $t_{\rm c}$. On the other hand, the median Lagrangian protohalo AM is $J^{\rm L}(M,t_{\rm c})/2$.
Both the mean and median Lagrangian AMs are independent of the (arbitrary) initial time $t_{\rm i}$ and of the cosmic time $t$, as expected, while they depend on the protohalo collapse time $t_{\rm c}$ through $\delta(t_{\rm c},t_{\rm i})/D(t_{\rm i})$ and on the halo virial mass $M$ essentially through factor $M^{5/3}$, in agreement with the results of simulations (e.g. Sugerman, Summers & Kamionkowski 2000). In turn, their Eulerian counterparts depend $t$ through the additional factor $a^{2}(t)\dot{D}(t)$.
No need to say that the Lagrangian protohalo AM derived here only holds for objects in linear regime. Moreover, it presumes the protohalo is linear at all relevant smaller scales, implying that they are roughly homogeneous with inner local density equal, to leading order, to the mean cosmic density. In Paper II we will extend those results to protohaloes endowed with their real typical density profile (Salvador-Solé et al., 2012a) and infer the typical spin and specific AM profile of virialised haloes.
@@ Appendix A Tidal Tensor
${\bf T}_{{\rm d}ij}\!\equiv\!\frac{a^{3}(t_{\rm i})}{D(t_{\rm i})}\frac{\partial^{2}\Phi_{\rm d}}{\partial x_{j}\partial x_{i}}\!=\!\frac{a^{3}(t_{\rm i})}{D(t_{\rm i})}\frac{4\pi\Delta[S({\bf 0})]}{\Delta[S({\bf x})]}G\delta_{\rm d}\bar{\rho}(t_{\rm i})\!\!\left(\!\frac{x_{i}x_{j}}{[a_{{\rm d}i}^{2}\!+\!S({\bf x})][a_{{\rm d}j}^{2}\!+\!S({\bf x})]}\!\left\{\sum_{k=1}^{3}\!\frac{x_{k}^{2}}{[a_{{\rm d}k}^{2}+S(\bf x)]^{2}}\!\right\}^{\!\!-1}\!\!\!\!-\delta_{ij}\!\int_{S({\bf x})}^{\infty}\!\!\frac{\Delta[S({\bf x})]\,{\rm d}s}{\Delta(s)(a^{2}_{{\rm d}i}\!+\!s)}\!\right)\!,$ (50)
@@
Since $R_{\rm d}$ is limited by the condition $R_{\rm d}^{\rm th}=r$, the c.o.m. of the protohalo is at the edge of the ellipsoidal mass excess or default, i.e. $\sum_{k=1}^{3}x_{k}^{2}/a^{2}_{{\rm d}k}=1$ and $S({\bf x})=0$. Consequently, the tidal tensor is simply
${\bf T}_{{\rm d}ij}=\frac{a^{3}(t_{\rm i})}{D(t_{\rm i})}4\pi G\delta_{\rm d}\bar{\rho}(t_{\rm i})\left[\frac{x_{i}x_{j}}{a^{2}_{{\rm d}i}a^{2}_{{\rm d}j}}\left(\sum_{k=1}^{3}\frac{x_{k}^{2}}{a_{{\rm d}k}^{4}}\right)^{\!-1}-\delta_{ij}U(a^{2}_{{\rm d}k})\right]\!,$ (51)
@@
$U(a^{2}_{{\rm d}k})\equiv\frac{1}{2}\int_{0}^{\infty}\frac{{\rm d}s/a^{2}_{{\rm d}j}}{1+s/a_{{\rm d}j}^{2}}\left[\prod_{k=1}^{3}\left(1+\frac{s}{a^{2}_{{\rm d}k}}\right)\right]^{-1/2}=\frac{1}{2}\int_{1}^{\infty}\,\frac{{\rm d}y}{y^{3/2}}\left[\sum_{k=1}^{3}\frac{a_{{\rm d}j}+s(y)^{2}}{a_{{\rm d}k}^{2}+s(y)}\right]^{-1}.$ (52)
Taking into account that the ellipticity and prolateness of peaks is moderate (BBKS), we have
$U(a^{2}_{{\rm d}k})\approx\frac{1}{6}\int_{1}^{\infty}\,\frac{{\rm d}y}{y^{3/2}}=\frac{1}{3}.$ (53)
In addition, we can neglect the terms in $1-a^{2}_{{\rm d}k}/a^{2}_{\rm d}$, where $a_{\rm d}^{2}\equiv 1/3\sum_{k=1}^{3}a^{2}_{{\rm d}k}$, in the Taylor expansion of $\sum_{k=1}^{3}x_{k}^{2}/a^{4}_{{\rm d}k}$ around $\sum_{k=1}^{3}x_{k}^{2}/(a^{2}_{{\rm d}k}a^{2}_{\rm d})$,
$\sum_{k=1}^{3}\frac{x_{k}^{2}}{a_{{\rm d}k}^{4}}=\sum_{k=1}^{3}\frac{x_{k}^{2}}{a_{{\rm d}k}^{2}a_{\rm d}^{2}}\left[1+\left(1-\frac{a_{{\rm d}k}^{2}}{a^{2}_{\rm d}}\right)+...\right]\approx\frac{1}{a_{\rm d}^{2}}\sum_{k=1}^{3}\frac{x_{k}^{2}}{a_{{\rm d}k}^{2}}=\frac{1}{a_{\rm d}^{2}}.$ (54)
Thus, we arrive at the almost fully accurate expression
${\bf T}_{{\rm d}ij}\approx\frac{a^{3}(t_{\rm i})}{D(t_{\rm i})}\frac{4\pi}{3}G\delta_{\rm d}\bar{\rho}(t_{\rm i})\left[3\frac{x_{i}x_{j}}{a^{2}_{{\rm d}i}a^{2}_{{\rm d}j}}a_{\rm d}^{2}-\delta_{ij}\right].$ (55)
Note that in the limit $a_{{\rm d}k}=R^{\rm th}_{\rm d}$, where the two approximations above hold exactly, expression (55) recovers the Hessian of the gravitational potential at the edge ($r=R_{\rm d}^{\rm th}$) of a spherically symmetric system, as expected.